OpenStudy (anonymous):

collect like terms 7/3x + 5/6y - 2/3x - 1/2y + 9

OpenStudy (anonymous):

(7/3x)-(3/3x)=4/3x

OpenStudy (anonymous):

5/6y-3/2y=5/6y-9/6y

OpenStudy (anonymous):

=-4y

OpenStudy (anonymous):

soz not 4y :S

OpenStudy (anonymous):

4/3x+(2/6y=)1/3y+9

OpenStudy (anonymous):

(7/3x - 2/3x) + (5/6y-1/2y) + 9 (5/3x) + (5/6y - 3/6y) + 9 5/3x + 2/6y + 9 5/3x + 1/3 y + 9

OpenStudy (anonymous):

Your second x term is incorrect, MissM.

OpenStudy (dumbcow):

peddani is correct

OpenStudy (anonymous):

but 7-3+=4 not 5 :S

OpenStudy (anonymous):

but it's 7-2, not 7-3 :)

OpenStudy (anonymous):

oh i see i read the qs wrong :(

OpenStudy (anonymous):

I need help with 3 more . Collect like terms 26x + 8 - 16x -6

OpenStudy (anonymous):

Collect like terms 7p - p

OpenStudy (anonymous):

Do you know what it means to collect like terms?

OpenStudy (anonymous):

Yes you combine them i am just really bad at this

OpenStudy (anonymous):

Collect like terms 2x +11x

OpenStudy (anonymous):

Okay. Well, the first...let's rearrange it so it's easier to see. 26x - 16x + 8 - 6 Now, since 26 and 16 are both multiplied by x, you can add them together. And I assume you know how to add 8 and -6!

OpenStudy (anonymous):

you just have to remember to carry the signs with the coefficients. Coefficient just means the number multiplied by the variable - from the first question, 26 and -16 are the coefficients of x.

OpenStudy (anonymous):

alright i got that

OpenStudy (anonymous):

so what did you get for 26x + 8 - 16x - 6?

OpenStudy (anonymous):

wait so i add 26x + (-16x)

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

that is 0

OpenStudy (anonymous):

It is?

OpenStudy (anonymous):

oh unless the negative becomes a positive

OpenStudy (anonymous):

Is it? drop the x, for now -> 26 + (-16)=?

OpenStudy (anonymous):

it is 10

OpenStudy (anonymous):

Yes. so you are left with 10x + 8 - 6 = ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

8-6 = 2, yes, but you need to keep the 10x. So the final answer will be 10x - 2

OpenStudy (anonymous):

SORRY! 10x + 2

OpenStudy (anonymous):

ohhhhh because you need to first collect the like terms! for instance this is like proportions

OpenStudy (anonymous):

So, try your next one.

OpenStudy (anonymous):

7p - p... Keep in mind this is the same as 7p - 1p

OpenStudy (anonymous):

i know any letter "number" in algebra by its self is always 1

OpenStudy (anonymous):

Okay...so what is the simplification of this problem?

OpenStudy (anonymous):

the answer will be 6p?

OpenStudy (anonymous):

Yes! What about your last one? 2x +11x ?

OpenStudy (anonymous):

oh k its 13x

OpenStudy (anonymous):

You got it! :)

OpenStudy (anonymous):

thank you so much help

OpenStudy (anonymous):

You are welcome. You can give me a medal as thanks! :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

: ) ill prob will post a another question if i run into trouble with my homework sheet so keep an out out for my user name!

OpenStudy (anonymous):

I will, if I am here. Good luck!

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

oh look im back again well now im on the next step which is Solve using the additions principle

OpenStudy (anonymous):

and the problem is x + 17 = 6